A boxcar traveling at 10 m/s approaches a string of 2 identical boxcars sitting stationary on the track. The moving boxcar collides and links with the stationary cars and they all move off together along the track. What is the final speed of the cars immediately after the collision? (You may take the mass of each boxcar to be 18,537 kg.)

m/s
Answer #1

let m = 18537 kg

vi = 10 ms
vf = ?

Apply conservation of momentum

final momentum = initial momentum

3*m*vf = m*vi

vf = vi/3

= 10/3

= 3.33 m/s